OpenStudy (anonymous):

solve for x. x-1/12 = 7/2

OpenStudy (anonymous):

\[x-\frac{ 1 }{ 2 }=\frac{ 7 }{ ?2 }\]

OpenStudy (anonymous):

ignore that "?" XD

OpenStudy (anonymous):

I still don't understand what "x" is... Thanks though!

OpenStudy (anonymous):

x is a variable

OpenStudy (anonymous):

you have to see what x =

OpenStudy (anonymous):

first we add 1/2 to both sides

OpenStudy (anonymous):

soo what 1/2 - 1/2? = 0 right so one the left side of the equal sign we have just x x=?

OpenStudy (anonymous):

sorry my bsd1/12

OpenStudy (anonymous):

x=6?

OpenStudy (anonymous):

bad*

OpenStudy (anonymous):

x= 1/12 +7/2 thats the equation now

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x=\frac{ 43 }{ 12 }\]

OpenStudy (anonymous):

we added 1/12 to both sides now 1/12 + 7/2 = 43/12

OpenStudy (anonymous):

so x=43?

OpenStudy (anonymous):

do you know how we got 43/ 12?

OpenStudy (anonymous):

no 43/12 is the awnser x = 43/12